OpenStudy (anonymous):

HELP ME 5ln (x-1) =20

OpenStudy (anonymous):

that 5 can be put up as the exponent... (ln(x-1))^5 = 20 exponentiate.. e^(ln(1-x))^5 = e^20 (1-x)^5 = e^20 fifth root: \[1-x = \sqrt[5]{e ^{20}}\] subtract 1 from both sides and multiply by -1: \[x = -\sqrt[5]{e ^{20}} +1\]

OpenStudy (ranga):

Divide both sides by 5: ln(x-1) = 4 Raise both sides to the power e. On the left e and ln will cancel out leaving with just (x-1) x - 1 = e^4 add 1 to both sides. x = e^4 + 1

OpenStudy (anonymous):

so which on is correct?

OpenStudy (anonymous):

one***

OpenStudy (anonymous):

they are both right except I accidentally switched x-1 to 1-x in the middle of the problem. So, my answer is actually: \[x = \sqrt[5]{e ^{20}} + 1\] Both give the same result though